OpenStudy (anonymous):

Peeps did I even do this right? I needed to write this as a single fraction: 3 + (x/5) So, 3 + (x/5) = 3(5/5) + x/5 = 15 + x / 5 Is that correct? Or? If not, can you help me out por favor :/

OpenStudy (anonymous):

\[3 + \frac{x}{5} = 3*\frac{5}{5} + \frac{x}{5} = \frac{15}{5} + \frac{x}{5} = \frac{x + 15}{5}\]

OpenStudy (anonymous):

Silly you, how could 3 + x/5 be equal to 15 + x/5. :)

OpenStudy (anonymous):

Ok got it !